# Trade-offs in Data Memorization via Strong Data Processing Inequalities

### v1

@@ 4.3 Sparse Boolean Hypercube
In the sparse Boolean hypercube clustering problem $\mathbf{P}_{\mathrm{sB}}$, assume $\nu=Cd^{-1/2}$ for some sufficiently large absolute constant $C>0$, and that $n\leq c\log d$ for some sufficiently small absolute constant $c>0$. Then the following hold:
@@
  • Any algorithm $\mathcal{A}$ with $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})\leq\alpha<\frac{1}{2}$ satisfies
    $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})=\Omega\left(\frac{(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)}{\nu^{2}2^{2n}}\right)-O(\sqrt{d}\log d)=\Omega\left(\frac{d}{2^{2n}}\right).$
@@
The proof of Theorem 4.4 appears in Appendix D. To prove the memorization lower bound, we establish a test/train $\rho_{n}$-SDPI for $\rho_{n}\approx\nu^{2}2^{2n}$. To do so, we introduce a dominating variable $Z_{\theta}^{\mathrm{train}}\sim\mathrm{Bin}_{\frac{1}{2}+\xi}(X)$ for $X\sim\mathcal{P}_{\theta},~{}\xi\approx\nu 2^{n}$, and show that processing it into $\tilde{X}_{1:n}$ by fixing each coordinate with some suitable probability, or else drawing each sample independently along that coordinate, results in a training set which is identically distributed as $X_{1:n}$. Then, we avoid the need of a data generation SDPI by directly upper bounding the entropy of $\theta$ by $H(\theta)\lesssim\sqrt{d}\log d$, hence obtaining by the chain rule $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})\geq I(\mathcal{A}(X_{1:n});X_{1:n})-H(\theta)\gtrsim\frac{1}{\rho_{n}}-\sqrt{d}\log d=\Omega\left(\frac{d}{2^{2n}}\right)$, the latter holding under our assumptions on $\nu$ and $n$. The nearly-matching upper bound follows by considering an algorithm which only stores a subset of $O(d/2^{2n})$ coordinates in which the sample is constant, and arguing that sufficiently many of them are indeed in $S$ with high probability, which suffices for generalization via standard concentration bounds.
@@ 5 Lower Bounds for Mixtures of Clusters
As in (Feldman, 2020; Brown et al., 2021), we will only consider product priors $\Pi$ in which frequencies of clusters are chosen independently up to a normalization constant. Specifically for a distribution $p$ over $[0,1]$, the product prior $\Pi_{p}^{k}$ is defined by independently sampling $p_{1},\dots,p_{k}\sim p^{k}$, and defining $\pi_{i}=\frac{p_{i}}{\sum_{i^{\prime}}p_{i^{\prime}}}$.
@@ 6 Discussion
All in all, our proof techniques and lower bounds for specific problem instances support the following intuition for the phenomenon of data memorization observed in practice: The tail of real-world data distributions contains many subproblems for which relatively little data is available. Non-trivial accuracy on these subproblems can be achieved by exploiting all the relatively weak correlations present between points in the dataset and unseen points from the same subproblem. This however, requires memorizing (almost) all the features of the available data points. More data allows the learning algorithm to average out some of the inherent randomness (or “noise”) in the features of the given examples, and thus increase the correlations with the features of an unseen point. In turn, this allows the learning algorithm to memorize fewer features of the training data on that subpopulation, specifically those with the strongest correlations.
Our work leaves open directions for future work. First, we believe our general framework can be applied to other problem instances, and the reduction is aimed at making our results of general use. In particular, it is interesting to understand whether there are problems which exhibit an even slower memorization decay than those we exhibit here. A related technical question is how excess memorization decays for the Gaussian clustering problem $\mathbf{P}_{\mathrm{G}}$ when $n\gtrsim\sqrt{d}$. We note that while our memorization lower bound still holds in this regime, we are not aware of a matching upper bound.
@@ Acknowledgments
GK is supported through an Azrieli Foundation graduate fellowship. XL is supported by a Google fellowship.
@@ Appendix A Proofs for Section 3
Given the lemmas above, we are ready to complete the proof of Proposition 3.5. Let $\tilde{B}$ be the approximation of $B$ which is given by the SDPI assumption, and denote $M:=\mathcal{A}(B),~{}\tilde{M}:=\mathcal{A}(\tilde{B})$. Applying Lemma A.2 and the $\rho$-SDPI assumption that holds for $(A,\tilde{B})$, we see that
@@ B.1 Gaussian sample complexity $(n=1)$
has error at most $0.01$. This will follow from standard Gaussian concentration bounds, since with high probability:
$X\sim\mathcal{P}_{\theta}~{}\implies~{}$ $\langle X,X_{1}\rangle\gtrsim\lambda^{2}d\gtrsim\sqrt{d}~{},$
$X\sim\mathcal{P}_{0}~{}\implies~{}$ $\mathop{\mathbb{E}}\langle X,X_{1}\rangle=0~{},\text{~{}~{}and hence~{}~{}}\left|\langle X,X_{1}\rangle\right|\lesssim\sqrt{d}~{}.$
Formally, start by noting that in the null case, $X_{1}$ and $X\sim\mathcal{P}_{0}$ are simply two independent isotropic Gaussians, and therefore a standard bound on their inner product ensures that
@@
Since $11-\lambda^{2}<$ 1, applying the same argument as in the null case to the latter three summands and union bounding ensures that
@@
Furthermore, a standard bound on the norm of a Gaussian vector (cf. Vershynin, 2018, Theorem 3.1.1) ensures that $\Pr[\|\theta\|^{2}<\gamma d]<\frac{1}{400}$ for some absolute constant $\gamma>0$. Union bounding over this event as well, we overall see that with probability at least $1-\frac{1}{200}:$
$\langle X,X_{1}\rangle\geq C^{2}\cdot\gamma d^{1/2}-2C\sqrt{4\log(1200)}d^{1/4}-\sqrt{4\log(1200)d}~{}.$
@@ B.2 Gaussian memorization lower bound
$Z_{\theta}^{\mathrm{train}}$ $=\mathcal{N}\left(\frac{\lambda\sqrt{n}}{\sqrt{1+(n-1)\lambda^{2}}}\cdot\theta,\frac{1-\lambda^{2}}{1+(n-1)\lambda^{2}}\cdot I_{d}\right),$ (2)
$Z_{\theta}^{\mathrm{test}}$ $=\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d}).$
To see that these variables satisfy the $\rho_{n}$-SDPI, we will use Fact 3.2, and therefore need to show the variables are $\sqrt{\rho_{n}}$-correlated unit Gaussians, for suitable $\rho_{n}$. To that end, recalling that $\theta\sim\mathcal{N}(0_{d},I_{d})$ and therefore marginalized over $\theta$ it holds that
$Z_{\theta}^{\mathrm{train}}$ $=\frac{\lambda\sqrt{n}}{\sqrt{1+(n-1)\lambda^{2}}}\mathcal{N}(0_{d},I_{d})+\mathcal{N}\left(0_{d},\frac{1-\lambda^{2}}{1+(n-1)\lambda^{2}}\cdot I_{d}\right)$
$=\mathcal{N}\left(0_{d},\frac{\lambda^{2}n}{1+(n-1)\lambda^{2}}\cdot I_{d}\right)+\mathcal{N}\left(0_{d},\frac{1-\lambda^{2}}{1+(n-1)\lambda^{2}}\cdot I_{d}\right)$
$=\mathcal{N}(0_{d},I_{d})~{},$
and similarly
$Z_{\theta}^{\mathrm{test}}=\lambda\cdot\mathcal{N}(0_{d},I_{d})+\mathcal{N}(0_{d},(1-\lambda^{2})I_{d})=\mathcal{N}(0_{d},I_{d})~{}.$
We see that $Z^{\mathrm{train}}_{\theta},Z^{\mathrm{test}}_{\theta}$ are both unit Gaussians, with each coordinate $i\in[d]$ satisfying
@@
We turn to argue about the existence of the desired mappings $\Phi^{\mathrm{train}},\Phi^{\mathrm{test}}$. First note that $Z_{\theta}^{\mathrm{train}}$ can be mapped to $\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ simply by rescaling by $\sqrt{\frac{n}{1+(n-1)\lambda^{2}}}$. Furthermore, it is known that $\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ can be processed to produce $X_{1},\dots,X_{n}\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d})$ since the average $\frac{1}{n}\sum_{i=1}^{n}X_{i}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ is a sufficient statistic of Gaussians with the same known variance (cf. Cover and Thomas, 1999, Section 2.9). By composition, this provides $\Phi^{\mathrm{train}}$. Moreover, $\Phi^{\mathrm{test}}$ is simply the identity since $Z^{\mathrm{test}}_{\theta}\sim\mathcal{P}_{\theta}$. We therefore establish the test/train condition in Theorem 3.8 with $\delta_{n}\equiv 0$.
To establish the data generation SDPI with $\epsilon_{n}\equiv 0$ as well, we once again use our construction of the post-processing $Z^{\mathrm{train}}_{\theta}$ as in Eq. (2), and claim that $\theta$ and $Z^{\mathrm{train}}_{\theta}$ satisfy the $\tau_{n}$-SDPI for $\sqrt{\tau_{n}}=\frac{\lambda\sqrt{n}}{\sqrt{1+(n-1)\lambda^{2}}}$. This follows as we know that they are both unit Gaussians, and by noting that they are $\sqrt{\tau_{n}}$-correlated coordinate-wise by construction of $Z^{\mathrm{train}}_{\theta}$. Therefore, the dataset $X_{1:n}$ which we have shown to be a post-processing of $Z_{\theta}^{\mathrm{train}}$ also satisfies $\tau_{n}$-SDPI with respect to $\theta$, since for every $M$ such that $M\perp\!\!\!\perp\theta\,|\,X_{1:n}$ we can use the SDPI and the (regular) DPI to see that
$I(M;\theta)\leq\tau_{n}I(M;Z^{\mathrm{train}}_{\theta})\leq\tau_{n}I(M;X_{1:n}).$
Overall, by applying Theorem 3.8, we see that the conditions of Theorem 3.7 are met, and we get that for any algorithm $\mathcal{A}$ with $\mathrm{err}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\leq\alpha<\frac{1}{2}$
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{G}})\geq\frac{1-\tau_{n}}{\rho_{n}}C_{\alpha}=\frac{1-\frac{\lambda^{2}n}{1+(n-1)\lambda^{2}}}{\frac{\lambda^{4}n}{1+(n-1)\lambda^{2}}}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)=\frac{1-\lambda^{2}}{\lambda^{4}n}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~{}.$
@@ B.3 Gaussian memorization upper bound
Given a sample $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, we will show that there exists an algorithm that returns a good classifier with the claimed memorization upper bound. For that, we argue that an algorithm can return a high accuracy classifier $h=\mathcal{A}(X_{1:n})$ which is describable using $O(d\log(\tfrac{d}{n})/n)$ bits, hence in particular

### v2

@@ 4.3 Sparse Boolean Hypercube
In the sparse Boolean hypercube clustering problem $\mathbf{P}_{\mathrm{sB}}$, assume $\nu=C/\sqrt{d}$ for some sufficiently large absolute constant $C>0$ and $n\leq c\log d$ for a sufficiently small constant $c>0$. Then the following hold:
@@
  • For any constant $\alpha<1/2$,
    $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{sB}},\alpha)=\Omega\left(\frac{(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)}{\nu^{2}2^{2n}}\right)-O(\sqrt{d}\log d).$
    In particular, $\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{sB}})=\Omega\left(\frac{d}{2^{2n}}\right)$.
@@
The proof of Theorem 4.4 appears in Appendix D. To prove the memorization lower bound, we establish a test/train $\rho_{n}$-SDPI for $\rho_{n}\approx\nu^{2}2^{2n}$. To do so, we introduce a dominating variable $Z_{X}^{\mathrm{train}}\sim\mathrm{BSC}_{\frac{1}{2}+\xi}(X)$ for $X\sim\mathcal{P}_{\theta},~\xi\approx\nu 2^{n}$, and show that processing it into $\tilde{X}_{1:n}$ by fixing each coordinate with some suitable probability, or else drawing each sample independently along that coordinate, results in a training set which is distributed identically to $X_{1:n}$. We note that this is a more direct way to establish the test/train SDPI than the one we give in Theorem 3.8. We detail this approach in Lemma D.1 and Theorem D.2. Then, we directly upper-bound $I(\mathcal{A}(X_{1:n}),\theta)$ by the entropy of $\theta$ that satisfies $H(\theta)\lesssim\sqrt{d}\log d$. We then conclude that $\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})\geq I(\mathcal{A}(X_{1:n});X_{1:n})-H(\theta)\gtrsim\frac{1}{\rho_{n}}-\sqrt{d}\log d=\Omega\left(\frac{d}{2^{2n}}\right)$, the latter holding under our assumptions on $\nu$ and $n$. The nearly-matching upper bound follows by considering an algorithm which only stores a subset of $O(d/2^{2n})$ coordinates in which the sample is constant, and arguing that sufficiently many of them are indeed in $S$ with high probability, which suffices for generalization via standard concentration bounds.
@@ 5 Lower Bounds for Mixtures of Clusters
As in (Feldman, 2020; Brown et al., 2021), we will only consider product priors $\Pi$ in which frequencies of clusters are chosen independently up to a normalization constant. Specifically, for a distribution $p$ over $[0,1]$, the product prior $\Pi_{p}^{k}$ is defined by independently sampling $p_{1},\dots,p_{k}\sim p^{k}$, and defining $\pi_{i}=\frac{p_{i}}{\sum_{i^{\prime}}p_{i^{\prime}}}$.
@@ 6 Discussion
All in all, our proof techniques and lower bounds for specific problem instances support the following intuition regarding the phenomenon of data memorization observed in practice: The tail of real-world data distributions contains many subproblems for which relatively little data is available. Non-trivial accuracy on these subproblems can be achieved by exploiting all the relatively weak correlations present between points in the dataset and unseen points from the same subproblem. This, however, requires memorizing (almost) all the features of the available data points. More data allows the learning algorithm to average out some of the inherent randomness (or “noise”) in the features of the given examples, thus increasing the correlations with the features of an unseen point. In turn, this allows the learning algorithm to memorize fewer features of the training data for that subpopulation, specifically those with the strongest correlations.
Our work leaves open directions for future work. First, we believe our general framework can be applied to other problem instances, and the reduction is aimed at making our results of general use. In particular, it is interesting to understand whether there are problems that exhibit an even slower memorization decay than those we present here.
@@ Acknowledgments
We are grateful to Tor Lattimore (Lattimore, 2025) for permission to include his algorithm for Gaussian clustering in Section B.3 in the paper. GK is supported through an Azrieli Foundation graduate fellowship. XL is supported by a Google fellowship.
@@ Appendix A Proofs for Section 3
[…]
@@ B.1 Gaussian sample complexity $(n=1)$
has expected error at most $0.01$. This will follow from standard concentration bounds which we recall below (cf. Vershynin, 2018).
Lemma B.1.
Suppose $X,Y\overset{iid}{\sim}\mathcal{N}(0_{d},I_{d})$ are independent isotropic Gaussians. Then the following hold:
  1. 1.
    $\Pr[\langle X,Y\rangle\geq t]\leq\exp(-t^{2}/2d)$ for all $t>0$;
  2. 2.
    $\Pr[\|X\|^{2}\leq\gamma d]\leq\exp(-(1-\gamma)^{2}d/4)$ for all $\gamma\in(0,1)$.
We start by noting that in the null case, $X_{1}$ and $X\sim\mathcal{P}_{0}$ are simply two independent isotropic Gaussians, and therefore Lemma B.1.1 ensures that
@@
$=\frac{C^{2}}{d^{1/2}}\|\theta\|^{2}+\frac{C}{d^{1/4}}\langle\theta,g_{0}\rangle+\frac{C}{d^{1/4}}\langle\theta,g_{1}\rangle+\langle g_{0},g_{1}\rangle~.$ (2)
Since $1-\lambda^{2}<1$, applying Lemma B.1.1 to the latter three summands and union bounding ensures that
@@
Furthermore, Lemma B.1.2 ensures that for $d$ larger than some absolute constant, it holds that $\Pr[\|\theta\|^{2}<d/2]<\frac{1}{400}$ . Union bounding over this event as well, we see that with probability at least $1-\frac{1}{200}:$
$\langle X,X_{1}\rangle\geq\frac{1}{2}C^{2}d^{1/2}-2C\sqrt{4\log(1200)}d^{1/4}-\sqrt{4\log(1200)d}~.$
@@ B.2 Gaussian memorization lower bound
$Z_{\theta}^{\mathrm{train}}$ $\sim\mathcal{N}\left(\frac{\lambda\sqrt{n}}{\sqrt{1+(n-1)\lambda^{2}}}\cdot\theta,\frac{1-\lambda^{2}}{1+(n-1)\lambda^{2}}\cdot I_{d}\right),$ (3)
$Z_{\theta}^{\mathrm{test}}$ $\sim\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d}).$
To see that these variables satisfy the $\rho_{n}$-SDPI, we will use Fact 3.2, and therefore need to show the variables are $\sqrt{\rho_{n}}$-correlated jointly distributed unit Gaussians, for suitable $\rho_{n}$. To that end, we denote by $Z^{\mathrm{train}},~Z^{\mathrm{test}}$ the marginals of $Z_{\theta}^{\mathrm{train}},~Z_{\theta}^{\mathrm{test}}$ over $\theta\sim\mathcal{N}(0_{d},I_{d})$ respectively. Note that Eq. (3) shows that $Z_{\theta}^{\mathrm{train}},~Z_{\theta}^{\mathrm{test}}$ are jointly Gaussian whose marginals satisfy $Z^{\mathrm{train}},Z^{\mathrm{test}}\sim N(0_{d},I_{d})$ simply by summing means and variances. We further note that each coordinate $i\in[d]$ satisfies
@@
We turn to argue about the existence of the desired mappings $\Phi^{\mathrm{train}},\Phi^{\mathrm{test}}$. First note that $Z_{\theta}^{\mathrm{train}}$ can be mapped to a sample from $\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ simply by rescaling by $\sqrt{\frac{n}{1+(n-1)\lambda^{2}}}$. We further use a known result that states that the sample average $\frac{1}{n}\sum_{i=1}^{n}X_{i}\sim\mathcal{N}(\lambda\theta,\frac{1-\lambda^{2}}{n}I_{d})$ is a sufficient statistic for the mean of a Gaussian distribution with known covariance matrix (cf. Cover and Thomas, 1999, Section 2.9), rephrased below:
Lemma B.2.
[…]
A composition of scaling and the result above provides $\Phi^{\mathrm{train}}$. Moreover, $\Phi^{\mathrm{test}}$ is simply the identity since $Z^{\mathrm{test}}_{\theta}\sim\mathcal{P}_{\theta}$. We therefore establish the test/train condition in Theorem 3.8 with $\delta_{n}\equiv 0$.
To establish the data generation SDPI with $\epsilon_{n}\equiv 0$ as well, we once again use our construction of the post-processing $Z^{\mathrm{train}}_{\theta}$ as in Eq. (3), and claim that $\theta$ and $Z^{\mathrm{train}}_{\theta}$ satisfy the $\tau_{n}$-SDPI for $\sqrt{\tau_{n}}:=\frac{\lambda\sqrt{n}}{\sqrt{1+(n-1)\lambda^{2}}}$. This follows as we know that they are jointly unit Gaussians, and by noting that they are $\sqrt{\tau_{n}}$-correlated coordinate-wise by construction of $Z^{\mathrm{train}}_{\theta}$.
Overall, by applying Theorem 3.8, we see that the conditions of Theorem 3.7 are met, and we get that for any $\alpha<\frac{1}{2}$,
$\mathrm{mem}_{n}(\mathbf{P}_{\mathrm{G}},\alpha)\geq\frac{1-\tau_{n}}{\rho_{n}}C_{\alpha}=\frac{1-\frac{\lambda^{2}n}{1+(n-1)\lambda^{2}}}{\frac{\lambda^{4}n}{1+(n-1)\lambda^{2}}}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)=\frac{1-\lambda^{2}}{\lambda^{4}n}\cdot(1-2\alpha)\log\left(\frac{1-\alpha}{\alpha}\right)~.$
@@ B.3 Gaussian memorization upper bound
We now describe a matching upper bound for the task of Gaussian clustering. The first upper bound is based on adding Gaussian noise and matches the lower bound in the entire range of $n$. It was communicated to us by Tor Lattimore. For completeness we also include an algorithm based on a projection from an earlier version of this work. It gives an upper bound that is tight up to a logarithmic factor and applies only when $n\leq c\sqrt{d}$ for some constant $c>0$.
B.3.1 Gaussian noise addition
Setup and the algorithm.
Let $\lambda=Cd^{-1/4}$ where $C$ is some absolute constant. Recall that the goal is to distinguish $\mathcal{P}_{\theta}=\mathcal{N}(\lambda\theta,(1-\lambda^{2})I_{d})$ from $\mathcal{P}_{0}=\mathcal{N}(0_{d},I_{d})$ where $\theta\sim\mathcal{N}(0_{d},I_{d})$.
We describe the algorithm: on input $X_{1:n}$ the algorithm calculates $\hat{X}:=\frac{1}{n}\sum_{i\in[n]}X_{i}$ and $\tilde{\mu}=\hat{X}+\eta$ where $\eta\sim\mathcal{N}(0_{d},I_{d})$ is an independent Gaussian. Then it returns the linear threshold classifier $\mathcal{A}(X_{1:n})(X)=\mathbbm{1}\left\{\langle X,\tilde{\mu}\rangle\geq\tau\right\}$ where $\tau$ will be defined later.
Mutual information bound.
With the setup above, consider the joint random variable $(\theta,X_{1:n},\hat{X},\eta,\mathcal{A}(X_{1:n}))$. The classifier $\mathcal{A}(X_{1:n})$ is as a post-processing of $\hat{X}$ and $\eta$ where we recall that $\eta\sim\mathcal{N}(0_{d},I_{d})$ is an independent Gaussian. Thus,
$I(\mathcal{A}(X_{1:n});X_{1:n}\mid\theta)\leq I(\hat{X}+\eta;\hat{X}\mid\theta)=\mathop{\mathbb{E}}_{\theta,u\sim\hat{X}}\left[D_{\mathrm{KL}}\left(u+\eta\,||\,\hat{X}+\eta\right)\right]$
Note that conditioned on $\theta$ and $u\sim\hat{X}$, the random variable $u+\eta$ is distributed as $\mathcal{N}(u,I_{d})$ while the random variable $\hat{X}+\eta$ (with an independently drawn $\hat{X}$) is distributed as $\mathcal{N}(\theta,(\frac{1-\lambda^{2}}{n}+1)I_{d})$.
Recall that the KL divergence between from Gaussian $\mathcal{N}(\mu_{1},\Sigma_{1})$ to Gaussian $\mathcal{N}(\mu_{2},\Sigma_{2})$ is given by
$D_{\mathrm{KL}}\left(\mathcal{N}(\mu_{1},\Sigma_{1})\,||\,\mathcal{N}(\mu_{2},\Sigma_{2})\right)=\frac{1}{2}\left(\log{\frac{\det\Sigma_{2}}{\det\Sigma_{1}}}-d+\mathrm{Tr}(\Sigma_{2}^{-1}\Sigma_{1})+(\mu_{2}-\mu_{1})^{\top}\Sigma_{2}^{-1}(\mu_{2}-\mu_{1})\right).$
We specialize the formula to our setting, yielding
$I(\mathcal{A}(X_{1:n});X_{1:n}\mid\theta)$ $=\mathop{\mathbb{E}}_{\theta\sim\mathcal{N}(0,I_{d}),u\sim\mathcal{N}(\lambda\theta,\frac{(1-\lambda^{2})}{n}I_{d})}\left[D_{\mathrm{KL}}\left(\mathcal{N}(u,I_{d})\,||\,\mathcal{N}\left(\theta,\left(\frac{1-\lambda^{2}}{n}+1\right)I_{d}\right)\right)\right]$
$=\frac{1}{2}\mathop{\mathbb{E}}\left[d\log\frac{(1-\lambda^{2})/n+1}{1}-d+d\cdot\frac{1}{(1-\lambda^{2})/n+1}+\frac{\|u-\theta\|^{2}}{(1+(1-\lambda^{2})/n)}\right]$
$=\frac{1}{2}\left(d\log(1+O(1/n))-d+d\cdot(1-O(1/n))+O(d/n)\right)$
$=O\left(\frac{d}{n}\right).$
where the approximation hides absolute constants independent of $d$ and $\lambda$, and the inequality holds true for all sufficiently large $d$.
Accuracy of the algorithm.
We turn to establish the accuracy guarantee of our algorithm. Suppose $X\sim\mathcal{P}_{0}=\mathcal{N}(0_{d},I_{d})$. We see that $\langle X,\tilde{\mu}\rangle$ is distributed as $\mathcal{N}(0,\|\tilde{\mu}\|^{2})$. Recall that $\tilde{\mu}=\hat{X}+\eta$ is distributed as $\mathcal{N}(0_{d},(\lambda^{2}+\frac{1-\lambda^{2}}{n}+1)I_{d})$. Then, by using Lemma B.1, we know that with probability $1-\left(\frac{1}{nd}\right)^{10}$, it holds that $\langle X,\tilde{\mu}\rangle\leq O(C\sqrt{d})$ where $C$ is the (large enough) constant from our choice of $\lambda=Cd^{-1/4}$.
On the other hand, if $X\sim\mathcal{P}_{\theta}$ then both $X$ and $\tilde{\mu}$ can be seen as distributed as $X=\lambda\theta+g_{0},~\tilde{\mu}=\lambda\theta+g_{1}$, where $g_{0}\sim\mathcal{N}(0_{d},(1-\lambda^{2})I_{d}),\,g_{1}\sim\mathcal{N}(0_{d},(\frac{1-\lambda^{2}}{n}+1)I_{d})$ are independent of one another as well as of $\theta$. Hence,
$\langle X,\tilde{\mu}\rangle$ $=\lambda^{2}\|\theta\|^{2}+\lambda\langle\theta,g_{0}\rangle+\lambda\langle\theta,g_{1}\rangle+\langle g_{0},g_{1}\rangle$
The rest of the proof is therefore implied by the concentration given by Lemma B.1, similarly to the arguments following Eq. (2), which we repeat here briefly. Since we have $\|\theta\|_{2}^{2}\geq\frac{d}{2}$ with probability at least $1-\frac{1}{400}$, whereas the three latter inner product terms have magnitude $O(\sqrt{d})$ with the same high probability, it follows that $\langle X,\tilde{\mu}\rangle=\Omega(C^{2}\sqrt{d})$. For sufficiently large absolute constant $C$, this is clearly larger than $\langle X,\tilde{\mu}\rangle=O(C\sqrt{d})$ when $X\sim\mathcal{P}_{0}$. Hence, by setting the threshold $\tau$ in between (e.g., $\tau=C^{3/2}\sqrt{d}$) we get a linear classifier with small error.
B.3.2 Projection-based algorithms
Given a sample $X_{1:n}\sim\mathcal{P}_{\theta}^{n}$, we will show that there exists a projection-based algorithm that returns a good classifier with the claimed memorization upper bound, whenever $n=O(\sqrt{d})$. For that, we argue that an algorithm can return a high accuracy classifier $h=\mathcal{A}(X_{1:n})$ which is describable using $O(d\log(\tfrac{d}{n})/n)$ bits. In particular,